OpenStudy (anonymous):

Im having trouble with taking the derivative of this: ((x)/sqrt(x^2+1)

OpenStudy (anonymous):

denominator is easy enough, it will be \(x^2+1\) numerator is going to be a drag

OpenStudy (anonymous):

If I used the product rule, it would be \[(x)(x^2+1)^(-1/2)\] correct? Or would the x be a 1?

OpenStudy (anonymous):

quotient rule is best product rule will make even more of a mess it is really not that bad

OpenStudy (anonymous):

I mean, I got \[\frac{ 1 }{(x^2+1)^3/2 }\]

OpenStudy (anonymous):

numerator will be \(g'f-fg'\)with \[f(x)=x,f'(x)=1,g(x)=\sqrt{x^2+1},g'(x)=\frac{x}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

I don't know. I always feel as if Im doing them completely wrong haha.

OpenStudy (anonymous):

plug and chug you will get a compound fraction that you can take care of by multiplying top and bottom by \(\sqrt{x^2+1}\)

OpenStudy (anonymous):

damn i made a typo above numerator is \(gf'-fg'\)

OpenStudy (anonymous):

write it out numerator will be \(\sqrt{x^2+1}\times 1-x\times \frac{x}{\sqrt{x^2+1}}\)

OpenStudy (anonymous):

or \[\sqrt{x^2+1}-\frac{x^2}{\sqrt{x^2+1}}\]

OpenStudy (anonymous):

denominator is \(x^2+1\)

OpenStudy (anonymous):

so you get a nice ugly compound fraction \[f'(x)=\frac{\sqrt{x^2+1}-\frac{x^2}{\sqrt{x^2+1}}}{x^2+1}\]

OpenStudy (anonymous):

then multiply top and bottom by \(\sqrt{x^2+1}\) and i guess you get what you wrote above

OpenStudy (anonymous):

\[\frac{1}{(x^2+3)^{\frac{3}{2}}}\]

OpenStudy (anonymous):

Whoa! Yeah, I wasn't sure at all though. haha Thanks a ton! :D

OpenStudy (anonymous):

Lol. I hope you realize how much that has helped me. lol Im going through these like nothing. Thanks!